OpenStudy (rmpjingwei):

What is the minimum distance between any point on the circle x square+y square=25 and the line y=−3/4x+75/4?

OpenStudy (mandre):

We will use a line perpendicular to line y=−3/4x+75/4. As we are working with a circle with it's center at the origin the line will go through the origin. That will be y=4/3x. If you multiply there slopes you get -1 which is how you determine 2 perpendicular lines. You can now calculate where they intersect by using the 2 equations: y = 4/3 x y= −3/4x+75/4 That means 4/3 x = −3/4x+75/4. i.e. x = 9. If you plug that into the first equation you get y = 4/3 (9) = 12. Using pythagoras the distance from the origin to the intersection of teh 2 lines are the square root of (9 ^2 + 12^2 ) = 15. So the distance from the circle is 15-radius = 15-5 = 10.